40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Entry  Wed Jul 15 22:23:17 2015, Koji, Electronics, AM Stabilized EOM Driver, E1400445 first look IMG_20150714_195534852.jpgIMG_20150714_195227746_HDR.jpg
    Reply  Sat Jul 18 11:37:21 2015, Koji, Electronics, AM Stabilized EOM Driver, D0900848 power board ~ oscillation issue solved 6x
       Reply  Wed Jul 22 09:43:01 2015, Koji, Electronics, AM Stabilized EOM Driver, Power supply test of the EOM/AOM Driver 
          Reply  Wed Jul 22 10:15:14 2015, Koji, Electronics, AM Stabilized EOM Driver, RF test of the EOM/AOM Driver S1500117 
             Reply  Sat Jul 25 17:24:11 2015, Koji, Electronics, AM Stabilized EOM Driver, RF test of the EOM/AOM Driver S1500117 7x
                Reply  Tue Jul 28 18:36:50 2015, Koji, Electronics, AM Stabilized EOM Driver, RF test of the EOM/AOM Driver S1500117 IMG_20150727_214536773_HDR.jpgEOM_Driver_DAQ_TF_test.pdfEOM_Driver_Mon_PSD.pdf
Message ID: 230     Entry time: Tue Jul 28 18:36:50 2015     In reply to: 229
Author: Koji 
Type: Electronics 
Category: AM Stabilized EOM Driver 
Subject: RF test of the EOM/AOM Driver S1500117 

Final Test Result of S1500117: https://dcc.ligo.org/LIGO-S1500117


After some staring the schematic and checking some TFs, I found that the DAQ channels for MON2 have a mistake in the circuit.
Differential driver U14 and U15 of D0900848 are intended to have the gain of +10 and -10 for the pos and neg outputs.
However, the positive output has the gain of +1.

Daniel suggested to shift R66 and R68 by one pad, replace with them by ~5.5K and add a small wire from the now "in air" pad to
the GND near pad 4.

The actual modification can be seen on Attachment 1. The resulting gain was +10.1 as the resisters of 5.49k were used.

The resulting transfer function is found in the Attachment 2. ow the nominal magnitude is ~x20.

You may wonder why the transfer function for MON1 is noisy and lower at low freq (f<1kHz). This is because the input noise of the FFT analizer
contributed to the BNC MON1 signal. High frequency part dominated the RMS of the signal and the FFT analyzer could not have proper range
for the floor noise. The actual voltage noise comparison between the BNC and DAQ signals for MON1 and MON2 can be found in Attachment 3.
 

Attachment 1: IMG_20150727_214536773_HDR.jpg  741 kB  | Hide | Hide all
IMG_20150727_214536773_HDR.jpg
Attachment 2: EOM_Driver_DAQ_TF_test.pdf  769 kB  | Hide | Hide all
EOM_Driver_DAQ_TF_test.pdf
Attachment 3: EOM_Driver_Mon_PSD.pdf  879 kB  | Hide | Hide all
EOM_Driver_Mon_PSD.pdf
ELOG V3.1.3-